OpenStudy (anonymous):

What is the domain and range of the quadratic equation y = –3(x – 10)2 + 5? Answers A)Domain: All Real Numbers Range: y 10 B)Domain: All Real Numbers Range: y 10 C)Domain: All Real Numbers Range: y 5 D)Domain: All Real Numbers Range: y 5 which one is the answer A,B,D,C ?

OpenStudy (apoorvk):

you mean y^2??

hero (hero):

Is anyone helping you with these @lilmagda105 ?

OpenStudy (anonymous):

nope

hero (hero):

I wish I had the time to help you, but I don't

OpenStudy (anonymous):

i can help

OpenStudy (anonymous):

do you know why I needed the vertex?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

less then or greater than goes with the answers

OpenStudy (across):

I will only do the first one:\[f(x)=-3(x-10)^2+5\]has a domain of \(\mathbb{R}\) and a range of\[f'(x)=-6(x-10)\to-6(x-10)=0\implies x=10\to f(10)=5\]\[\implies(-\infty, 5].\]

OpenStudy (anonymous):

sorry @across is right (10, 5) is your vertex.

OpenStudy (anonymous):

okay so the domain is all real numbers & the range is y is greater then 5 correct ?

OpenStudy (across):

No.

OpenStudy (across):

The range is \(y\leq5\).

OpenStudy (anonymous):

ohh thats what i was trying to say lol how do you use the equation symbols ?

OpenStudy (across):

Enclose them in either \.(\) or \.[\], without the dot.

OpenStudy (across):

That one up there, for example, is \.(y\leq5\) (again, without the dot).

OpenStudy (anonymous):

okay thanks